Circular No. 8095

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003cc, 2003cd, 2003ce, 2003cf Further to IAUC 8090, W. M. Wood-Vasey, G. Aldering, and P. Nu-

gent report the discoveries of additional apparent supernovae on unfiltered NEAT images.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003cc Feb. 3.53 10^{h}57^{m}29.^{s}60 +23^{o}35′34′′.7 20.3 12′′.3 E
2003cd Feb. 12.38 10 12 02.33 +11 16 39.1 19.5 5′′ W, 4′′ S
2003ce Mar. 11.28 10 12 10.43 −30 13 58.1 19.5 10′′ E, 5′′ N
2003cf Feb. 22.38 10 37 43.07 + 7 04 54.1 19.5 0′′.7 W, 4′′.4 S

Additional magnitudes from NEAT images: SN 2003cc, 2002 Mar. 22.41 UT, [20.6; 2003 Jan. 19.42, [20.3, Feb. 28.54, 19.5; Mar. 8.42, 19.5; 9.37, 19.5. SN 2003cd, Feb. 3.46, [20.6; 22.24, 19.5; 25.55, 19.8; 26.52, 19.8; 27.55, 20.1; Mar. 6.44, 20.0. SN 2003ce, 2002 Jan. 13.41, [21.0; 2003 Mar. 14.20, 19.3. SN 2003cf, 2002 May 23.18, [20.5; 2003 Feb. 20.51, [19.5; 28.45, 18.7; Mar. 6.47, 19.0; 14.17, 19.2; 14.44, 19.5.

XTE J1807–294 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); M. Juda, Harvard-Smithsonian Center for Astrophysics, and J. H. Swank, GSFC, report that the Chandra X-ray Observatory observed XTE J1807–294 on Mar. 10.8 UT, and the best-fit position was determined to be α = 18^{h}06^{m}59^{s}.8, δ = −29^{o}24′30′′ (equinox 2000.0; uncertainty ∼ 1′′). On Mar. 13.9, the RXTE PCA measured a 2–10-keV x-ray flux of 20 mCrab. Pulsations are still detected, and an orbital Doppler modulation is clearly evident. The best-fit orbital period is 40.0741 ± 0.0005 min, confirming it as the shortest period of any of the four accreting millisecond pulsars. Optical observations are encouraged.

COMET C/2002 X5 (KUDO-FUJIKAWA) Visual m_{1} estimates: Feb. 16.04 UT, 5.3 (J. D. Shanklin, Stanley,

Falkland Is., 10×50 binoculars); 23.97, 7.2 (A. Amorim, Florianopolis, Brazil, 7×50 binoculars); Mar. 3.41, 7.9 (D. A. J. Seargent, Cowra, N.S.W., 25×100 binoculars); 9.42, 8.8 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector); 12.82, 9.4 (N. Biver, Pico-Veleta, Spain, 0.20-m reflector).

2003 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green